# Dimension-independent convergence rates of randomized nets using median-of-means

### v1

@@ Abstract
Recent advances in quasi-Monte Carlo integration demonstrate that the median of linearly scrambled digital net estimators achieves near-optimal convergence rates for high-dimensional integrals without requiring a priori knowledge of the integrand’s smoothness. Building on this framework, we prove that the median estimator attains dimension-independent convergence under tractability conditions characterized by low effective dimensionality, a property known as strong tractability in complexity theory. Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for strong tractability.
@@ 1 Introduction
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows indefinitely while the variance remains bounded—a common scenario in path integral approximations [14]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [20], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can behave as if they reside in lower-dimensional subspaces. The connection to weighted function spaces is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [8]. A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error as the primary metric. In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates with high probability, avoiding worst-case error analysis entirely. This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases. Within this framework, we improve the convergence rates from nearly $O(n^{-\alpha})$ in [8] to nearly $O(n^{-\alpha-1/2})$, where $\alpha$ denotes the smoothness parameter in their analysis (we note that $\alpha$ is defined differently in this paper). Furthermore, we relax the strong tractability conditions, as detailed in Remarks 3 and 7. Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve strong tractability under conditions similar to those for polynomial lattices in [8], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [8].
@@ 2.2 Walsh functions and error analysis for base-2 digital nets
The equality holds in the $L^{2}$-sense. Using this decomposition, [24] establishes a key error structure for QMC estimators:
@@ 2.3 Probabilistic error guarantee of the median
The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [23, Lemma 4] for the univariate case $s=1$). These properties allow us to compute the variance of $\hat{\mu}_{\infty}$ and construct probabilistic error bounds by Chebyshev’s inequality.
@@
This is an intermediate result proven in [24, Lemma 6]. ∎
@@
We choose $\mathcal{A}$ to be the event $\{Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K\}$ for a carefully chosen index set $K\subseteq\mathbb{N}^{s}_{*}$. It is shown in [22, Subsection 2.4] that
@@ 2.4 Walsh coefficients and function smoothness
Equation (6) comes from [26, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [26, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $[0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [22, Remark 1].
@@ 3 ANOVA decomposition and effective dimensionality
Let $\boldsymbol{s}(\boldsymbol{k})=\{j\in 1{:}s\mid k_{j}\neq 0\}$ denote the support of $\boldsymbol{k}=(k_{1},\dots,k_{s})$. From the Walsh decomposition (3),
@@
which follows from the fact that $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})$ depends only on $\{x_{j}\mid j\in\boldsymbol{s}(\boldsymbol{k})\}$ and
@@
where the supremum is over all partitions $\mathcal{P}$ of $[0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$. This generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [17] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [22, Section 4.1]).
@@
$V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big{(}I_{v}(f\rho_{v^{c}})\Big{)}.$
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [22, Section 3.2],
$\Bigl{(}V^{v}_{\lambda}(f)\Bigl{)}^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl{(}V^{(|v|)}_{\lambda}\Bigl{(}f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}\Bigr{)}^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$ (12)
@@
Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for $v\subseteq u\subseteq 1{:}s,u\neq\varnothing$ and $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ with $\alpha_{j}=\alpha\ \forall j\in v$. Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha_{j}$ if $j\in u\setminus v$ and $|\kappa_{j}|=\alpha+1$ if $j\in v$,
@@
$B(\boldsymbol{k}_{u},v)=\{\boldsymbol{k}^{\prime}\in\mathbb{N}^{s}_{*}\mid\boldsymbol{s}(\boldsymbol{k}^{\prime})=u,\kappa^{\prime}_{j}=\kappa_{j}\text{if }j\in u\setminus v,\lceil\kappa^{\prime}_{j}\rceil_{1:(\alpha+1)}=\kappa_{j}\text{ if }j\in v\}.$ (13)
@@
Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$. Because $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$ and $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$, we can apply Lemma 3 to get
@@
From the proof of [22, Theorem 3], we can map $v$ to $1{:}|v|$ and show
@@
$\|f\|_{u,\alpha,\lambda}=\sup_{\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}}\sup_{v\subseteq\{j\mid\alpha_{j}=\alpha\}}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u}),$ (15)
where $f_{u}$ is the ANOVA component of $f$ defined in equation (8). We set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
@@
All subsequent results hold equivalently for $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ replacing $\|f\|_{\alpha,\lambda}$ and $\tilde{\gamma}_{u}$ replacing $\gamma_{u}$. However, since $\|f\|_{\alpha,\lambda}$ and $\gamma_{u}$ are intrinsic to $f$, while $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ requires external weights $\tilde{\gamma}_{u}$ unused by our algorithm, we adopt the unweighted framework for simplicity.
@@ 4 Tractability under the completely random design
In this section, we show the median RQMC estimator $\hat{\mu}^{(r)}_{\infty}$ under the completely random design achieves a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate given $\|f\|_{\alpha,\lambda}<\infty$, with no dimension-dependence if further $\gamma_{u}$ satisfies equation (34).
@@
$\mathcal{K}_{u,v}(T)=\Big{\{}(k_{j},j\in u)\in\mathbb{N}^{|u|}\Big{|}\begin{matrix}|\kappa_{j}|\leqslant\alpha\ \forall j\in u\setminus v,|\kappa_{j}|=\alpha+1\ \forall j\in v,\\ (\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|\leqslant T\end{matrix}.\Big{\}}$
@@
The next lemma quantifies how $T$ controls the size of $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$.
@@
$C_{\alpha,\lambda,\theta}=\frac{4^{-\alpha-1}}{\alpha!(4^{(\alpha+\lambda)(1-\theta)}-1)(4^{1-\theta}-1)^{\alpha}}+\sum_{\alpha^{\prime}=0}^{\alpha}\frac{4^{-\alpha^{\prime}}}{(\alpha^{\prime})!(4^{1-\theta}-1)^{\alpha^{\prime}}}.$ (18)
@@
and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$. It follows that $[t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$. Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big{(}\sum_{N=1}^{\infty}\rho_{\theta}^{N}\Big{)}^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(\rho_{\theta}^{-1}-1)^{\alpha^{\prime}}}.$ (20)
For $\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})$, we let $\kappa^{+}=\lceil\kappa\rceil_{1{:}\alpha}$ and $\ell=\lceil\kappa\rceil_{\alpha+1}$. By the one-to-one correspondence between $\kappa^{+}$ and $\kappa^{+}-\ell=\{\ell^{\prime}-\ell\mid\ell^{\prime}\in\kappa\}\in\mathcal{S}_{\alpha}(\mathbb{N})$,
@@
While larger $T$ implies a tighter bound on $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$, it also increases the probability $Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K_{u}(T)$. The next lemma provides a bound on the cardinality of $K_{u}(T)$.
@@
$|K_{u}(T)|\leqslant 2^{T/(\alpha+\lambda)}\Big{(}\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}+B_{\alpha,\lambda}\Big{)}^{|u|},$
@@
$A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=0}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}}.$ (21)
@@
The conclusion is trivial when $T<0$, so we assume $T\geqslant 0$. As in the proof Lemma 4, for each $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ we write $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$ and $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ for $j\in v$. By equation (13), each $\boldsymbol{k}^{\prime}=(k^{\prime}_{1},\dots,k^{\prime}_{s})\in B(\boldsymbol{k}_{u},v)$ must have the same $\kappa^{\prime}_{j}$ for $j\in 1{:}s\setminus v$. For $j\in v$, however, $\kappa^{\prime}_{j}\setminus\kappa_{j}$ can be any subset of $1{:}(\ell_{j}-1)$ if $\ell_{j}\geqslant 2$ and $\kappa^{\prime}_{j}=\kappa_{j}$ if $\ell_{j}=1$. Thus, $|B(\boldsymbol{k}_{u},v)|=\prod_{j\in v}2^{\ell_{j}-1}$. Because $B(\boldsymbol{k}_{u},v)$ for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ are mutually disjoint,

### v2

@@ Abstract
Recent advances in quasi-Monte Carlo integration demonstrate that the median of linearly scrambled digital net estimators achieves near-optimal convergence rates for high-dimensional integrals without requiring a priori knowledge of the integrand’s smoothness. Building on this framework, we prove that the median estimator attains dimension-independent convergence under tractability conditions characterized by low effective dimensionality, a property known as strong tractability in complexity theory. Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for dimension-independent convergence.
@@ 1 Introduction
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [17]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [23], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions. The connection to weighted function spaces [28] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [9]. A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error over the unit ball in a weighted Sobolev space as the primary metric. In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates on a fixed integrand with high probability, avoiding worst-case error analysis entirely. This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [14, 15]. Within this framework, we improve the convergence rates from nearly $O(n^{-\alpha})$ in [9] to nearly $O(n^{-\alpha-1/2})$, where $\alpha$ denotes the smoothness parameter in their analysis (we note that $\alpha$ is defined differently in this paper). Furthermore, we relax the conditions for dimension-independent convergence (Remarks 3 and 9). Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve dimension-independent convergence under conditions comparable to those for polynomial lattices in [9], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [9].
@@ 2.2 Walsh functions and error analysis for base-2 digital nets
The equality holds in the $L^{2}$-sense. Using this decomposition, [25, Theorem 1] establishes a key error structure for QMC estimators:
@@ 2.3 Probabilistic error guarantee of the median
The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [26, Lemma 4] for the univariate case $s=1$). These properties allow us to compute the variance of $\hat{\mu}_{\infty}$ and construct probabilistic error bounds by Chebyshev’s inequality.
@@
This is an intermediate result proven in [27, Lemma 6]. ∎
@@
We choose $\mathcal{A}$ to be the event $\{Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K\}$ for a carefully chosen index set $K\subseteq\mathbb{N}^{s}_{*}$. It is shown in [25, Subsection 2.4] that
@@ 2.4 Walsh coefficients and function smoothness
Equation (6) comes from [30, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [30, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $[0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [25, Remark 1].
@@ 3 ANOVA decomposition and effective dimensionality
Let $\boldsymbol{s}(\boldsymbol{k})=\{j\in 1{:}s\mid k_{j}\neq 0\}$ denote the support of $\boldsymbol{k}=(k_{1},\dots,k_{s})$. From the Walsh decomposition (3), we know that
@@
because $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})$ depends only on $\{x_{j}\mid j\in\boldsymbol{s}(\boldsymbol{k})\}$ and
@@
where the supremum is over all partitions $\mathcal{P}$ of $[0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$. When $d=0$ and $f\in\mathbb{R}$, $V^{(0)}_{\lambda}(f)=|f|$ by convention.
The Vitali variation generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [20] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [25, Section 4.1] for an example).
@@
$V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big{(}I_{v}(f\rho_{v^{c}})\Big{)}.$ (12)
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [25, Section 3.2],
$\Bigl{(}V^{v}_{\lambda}(f)\Bigl{)}^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl{(}V^{(|v|)}_{\lambda}\Bigl{(}f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}\Bigr{)}^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$
@@
Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for nonempty $u\subseteq 1{:}s$, $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and a subset $v\subseteq\{j\mid\alpha_{j}=\alpha\}$. Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$,
@@
$B(\boldsymbol{k}_{u},v)=\{\boldsymbol{k}^{\prime}\in\mathbb{N}^{s}_{*}\mid\boldsymbol{s}(\boldsymbol{k}^{\prime})=u,\ \lceil\kappa^{\prime}_{j}\rceil_{1:(\alpha+1)}=\kappa_{j}\ \forall j\in v,\ \kappa^{\prime}_{j}=\kappa_{j}\ \forall j\in u\setminus v\}.$ (13)
@@
Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$. Because $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$, we can apply Lemma 3 to get
@@
From the proof of [25, Theorem 3], we can map $v$ to $1{:}|v|$ and show
@@
$\|f\|_{u,\alpha,\lambda}=\sup_{v\subseteq u}\sup_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u}),$ (15)
where $f_{u}$ is the ANOVA component of $f$ defined in equation (8). Notice that we add the restriction $\alpha_{j}>0$ for $j\in u\setminus v$ because $k_{j}\in\mathbb{N}$ implies $\alpha_{j}=|\kappa_{j}|>0$ in the statement of Lemma 4. In particular, $\|f\|_{u,\alpha,\lambda}=V^{u}_{\lambda}(f_{u})$ when $\alpha=0$. We further set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
@@
All subsequent results hold equivalently with $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ replacing $\|f\|_{\alpha,\lambda}$ and $\tilde{\gamma}_{u}$ replacing $\gamma_{u}$. However, since $\|f\|_{\alpha,\lambda}$ and $\gamma_{u}$ are intrinsic to $f$, while $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ requires external weights $\tilde{\gamma}_{u}$ unused by our algorithm, we adopt the unweighted framework for simplicity.
@@ 4 Tractability under the completely random design
In this section, we show the median RQMC estimator $\hat{\mu}^{(r)}_{\infty}$ under the completely random design achieves near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates given $\|f\|_{\alpha,\lambda}<\infty$, with no dimension-dependence if further $\gamma_{u}$ satisfies equation (34).
@@
$\mathcal{K}_{u,v}(T)=\Big{\{}(k_{j},j\in u)\in\mathbb{N}^{|u|}\Big{|}\begin{matrix}|\kappa_{j}|=\alpha+1\ \forall j\in v,\ |\kappa_{j}|\leqslant\alpha\ \forall j\in u\setminus v,\\ (\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|\leqslant T\end{matrix}\Big{\}}.$
@@
The next lemma quantifies how $T$ controls the size of $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ for $\boldsymbol{k}\notin K_{u}(T)$.
@@
$C_{\alpha,\lambda,\theta}=\frac{4^{-\alpha}}{\alpha!(4^{(\alpha+\lambda)(1-\theta)}-1)(4^{1-\theta}-1)^{\alpha}}+\sum_{\alpha^{\prime}=0}^{\alpha}\frac{4^{-\alpha^{\prime}+1}}{(\alpha^{\prime})!(4^{1-\theta}-1)^{\alpha^{\prime}}}.$ (18)
@@
and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [6] for background knowledge on generating functions). It follows that $[t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$. Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}t^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}.$ (20)
with $t=\rho_{\theta}$. For $\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})$, we let $\kappa^{+}=\lceil\kappa\rceil_{1{:}\alpha}$ and $\ell=\lceil\kappa\rceil_{\alpha+1}$. By the one-to-one correspondence between $\kappa^{+}$ and $\kappa^{+}-\ell=\{\ell^{\prime}-\ell\mid\ell^{\prime}\in\kappa\}\in\mathcal{S}_{\alpha}(\mathbb{N})$,
@@
While larger $T$ implies a tighter bound on $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$, it also increases the probability that $Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K_{u}(T)$. The next lemma provides a bound on the cardinality of $K_{u}(T)$.
@@
$|K_{u}(T)|\leqslant 2^{T/(\alpha+\lambda)}\Big{(}A_{\alpha,\lambda}\max\big{(}T/(\alpha+\lambda),1\big{)}+B_{\alpha,\lambda}\Big{)}^{|u|},$
@@
$A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=1}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}},$ (21)
with the sum from $1$ to $\alpha$ interpreted as 00 when $\alpha=0$.
@@
The conclusion is trivial when $T<0$, so we assume $T\geqslant 0$. As in the proof Lemma 4, for each $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ we write $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$ and $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ for $j\in v$. By equation (13), each $\boldsymbol{k}^{\prime}=(k^{\prime}_{1},\dots,k^{\prime}_{s})\in B(\boldsymbol{k}_{u},v)$ must have the same $\kappa^{\prime}_{j}$ for $j\in 1{:}s\setminus v$. For $j\in v$, however, $\kappa^{\prime}_{j}\setminus\kappa_{j}$ can be any subset of $1{:}(\ell_{j}-1)$ if $\ell_{j}\geqslant 2$ and $\kappa^{\prime}_{j}=\kappa_{j}$ if $\ell_{j}=1$. Thus, $|B(\boldsymbol{k}_{u},v)|=\prod_{j\in v}2^{\ell_{j}-1}$. Because $B(\boldsymbol{k}_{u},v)$ are mutually disjoint for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$,